OpenStudy (anonymous):

True or False:\[\Large\int f(x)dx = \lim_{\Delta x \rightarrow 0}\Delta x\sum_{n=0}^{x}f\left(\dfrac{n}{\Delta x}\right)\] If false, what's the correct one?

OpenStudy (anonymous):

That's false.

OpenStudy (anonymous):

maybe\[\Large\int_{0}^{x} f(x)dx = \lim_{\Delta x \rightarrow 0}\Delta x\sum_{n=0}^{x}f\left(n{\Delta x}\right)\]

OpenStudy (anonymous):

still false

OpenStudy (anonymous):

The limit of a sum definition of an indefinite integral is:\[\bf \int\limits_{a}^{b}f(x) \ dx=\lim_{n \rightarrow \infty }\sum_{i=0}^{n}f(x_i) \Delta x\]

OpenStudy (anonymous):

sorry i meant to say definite integral not indefinite.

OpenStudy (anonymous):

Basically what it means is that if the function f(x), has its interval (a,b) divided into n subintervals each of length delta x, and i'th value is picked with in each subinterval, then as the number of those subintervals approaches infinity, delta x approaches 0, and the approximation of the area under the curve gets more and more and more accurate to the point where its exactly the area under the curve.

terenzreignz (terenzreignz):

I didn't know the indefinite integral had a definition in terms of summations :3

OpenStudy (anonymous):

false